# Probing vector chirality in the early Universe

### v1

@@ IV Measuring Helical asymmetry
On the other hand, in the parity asymmetric cases, the subsampled initial spin field shows notably large helical asymmetry with scale dependence similar to the full initial spin field. The significance of the asymmetry exceeds $6\sigma$ in the range of scale $0.3\leq k/[h\,{\rm Mpc}^{-1}]\leq 0.7$. For the halo spin field, we again detect considerable helical asymmetry. In particular, the positive departure from the perfect symmetry is detected more than $3\sigma$-significance on scales $0.35\leq k/[h\,{\rm Mpc}^{-1}]\leq 0.75$ and reaches about $5.1\sigma$ at maximum. On these scales, the halo spin field preserves on average more than $50\%$ of the initial helical asymmetry. This clearly demonstrates that the signature of vectorial-PV largely remains even after the gravitational collapse, persisting prominently across these non-linear scales.
V Conclusion and discussion
@@
We assessed the detectability of parity violation in the primordial vector fossil using galaxy spins. Utilizing simulations, we examined how effectively the primordial helical asymmetry is preserved in the late-time galaxy spin field. To this end, we developed a new method to generate initial conditions with parity violating four-point asymmetry as quantified by the initial spin field. Our approach provides a clean channel that effectively embeds parity asymmetry without altering the power spectrum or generating kurtosis via helical remapping, similarly in CMB lensing. This differentiates our approach from the implementation proposed in [50], where larger parity asymmetry inevitably accompanies larger kurtosis and variance.
@@
In the parity asymmetric simulations with $\sigma_{\mathcal{R}}=1.45$, we detected substantial helical asymmetry in the initial spin field, reaching a maximum amplitude of $\chi\simeq 0.18$. More than $50\%$ of such asymmetry on a range of scale survived in the late-time halo spin field and was detected maximally at $5.1\sigma$ significance. This demonstrates that the imprint of primordial vectorial-PV persists prominently in halo spins, only partly reduced by the non-linear gravitational evolution. On the other hand, in the null test (i.e., $\sigma_{\mathcal{R}}=0$), the initial spin was helically symmetric, and so was the halo spin field. Our findings validate the robustness of the halo spin field as tracers of the initial spin field, thereby highlighting its efficacy as a tool for probing the vectorial-PV in the early universe.
@@
Considering the upcoming Y1-data release of the Dark Energy Spectroscopic Instrument (DESI) [52], it presents a timely and compelling opportunity to investigate helical asymmetry in low-redshift galaxy spin fields. The DESI Bright Galaxy Sample [53] will provide high-quality photometric data for over 10 million nearby galaxies, offering precise measurements of position angle, axial ratio, color gradient [54] and winding direction [55], with which 3-dimensional galaxy spins can be unambiguously determined [32]. With such $10^{7}$ galaxy spins, one can potentially reduce the statistical uncertainty in helical asymmetry measurement by a factor of $1/\sqrt{10}$, compared to the present analysis using $10^{6}$ spins. Then, the current parity-violating model with $\sigma_{\mathcal{R}}=1.45$ can be detected up to $16\sigma$ significance in real observation.
@@
VI acknowledgments
We thank Donghui Jeong, Eiichiro Komatsu, Oliver Philcox, Shi-Fan Chen, and Tomomi Sunayama for providing valuable feedback on the manuscript. We also thank Academia Sinica Institute of Astronomy and Astrophysics (ASIAA) for hosting the workshop Large-scale Parity Violation Workshop during which this project was advanced. J.S. acknowledges the support by ASIAA. H.R.Y. is supported by National Science Foundation of China grant No. 12173030. T.O. acknowledges support from the Ministry of Science and Technology of Taiwan under grants No. MOST 111-2112-M- 001-061- and No. NSTC 112-2112-M-001-034- and the Career Development Award, Academia Sinica (AS-CDA-108-M02) for the period of 2019-2023. We thank ASIAA for providing computing resources (High Performance Computing Systems).

### v2

@@ Abstract
On the other hand, in the parity-violating realizations, the subsampled initial spin field shows notably large helical asymmetry, with scale dependence resembling that of the full initial spin field. The significance of the signal exceeds $6\sigma$ in the range of scale $0.3\leq k/[h\,{\rm Mpc}^{-1}]\leq 0.7$. For the halo spin field, we again detect considerable helical asymmetry. In particular, the positive departure from perfect symmetry is detected more than $3\sigma$-significance on scales $0.35\leq k/[h\,{\rm Mpc}^{-1}]\leq 0.75$, reaching a maximum of about $5.1\sigma$. On these scales, the halo spin field preserves on average more than $50\%$ of the initial helical asymmetry. This clearly demonstrates that the signature of vectorial-PV largely remains even after the gravitational collapse, persisting prominently across these non-linear scales.
@@
Conclusion and discussion.— We assessed the detectability of parity violation in primordial vector fossils using galaxy spins. Utilizing simulations, we examined how effectively the primordial helical asymmetry is preserved in the late-time galaxy spin field. To this end, we developed a new method to generate initial conditions with parity-violating four-point asymmetry as quantified by the power spectrum of the quadratic initial spin field. Our approach provides a clean channel that effectively embeds parity asymmetry without altering the power spectrum or generating kurtosis via helical remapping, similarly to CMB lensing. This differentiates our approach from the implementation proposed in [56], where larger parity asymmetry inevitably accompanies larger kurtosis and variance.
@@
In the parity asymmetric simulations with $\sigma_{\mathcal{R}}=1.45$, we detected substantial helical asymmetry in the initial spin field, reaching a maximum amplitude of $\chi\simeq 0.18$. More than $50\%$ of such asymmetry on a range of scale survived in the late-time halo spin field and was detected maximally at $5.1\sigma$ significance from a total volume of $0.5h^{-3}{\rm Gpc^{3}}$. This demonstrates that the imprint of primordial vectorial-PV persists prominently in halo spins, only partly reduced by the non-linear gravitational evolution. On the other hand, in the null test (i.e., $\sigma_{\mathcal{R}}=0$), the initial spin was helically symmetric, and so was the halo spin field. Our findings validate the robustness of the halo spin field as tracers of the initial spin field, thereby highlighting its efficacy as a tool for probing the vectorial-PV in the early universe.
@@
Considering the upcoming Y1-data release of the Dark Energy Spectroscopic Instrument (DESI) [58], it presents a timely and compelling opportunity to investigate helical asymmetry in low-redshift galaxy spin fields. The DESI Bright Galaxy Sample [59] will provide high-quality photometric data for over 10 million nearby galaxies, offering precise measurements of position angle, axial ratio, color gradient [60], projected rotational velocity, and winding direction [61], with which 3-dimensional galaxy spins can be unambiguously determined [35].
In light of this, we perform a simple forecast on the detection significance in observation. To this end, we create mock observed galaxy spin fields from the halo spin fields by reorienting halo spin vectors twice by $\theta$ in random directions. The two reorientations reflect that galaxy spins are typically more misaligned from their initial spins than halo spins [37] and that spin determination is subject to observational errors. For our median halo mass, the average angle between a halo and its galaxy spins is $21^{\circ}$ [37]. Thus, we adopt $\theta=21^{\circ}$ and assume the observational error is identical to $\theta$ for simplicity. In the mock galaxy spin fields, the helical asymmetry decreases by $18\%$ in $0.35<k/[h{\rm Mpc}^{-1}]<0.75$ relative to the halo spin fields. With such $10^{7}$ galaxy spins, the maximum detection significance for our parity-violating model with $\sigma_{\mathcal{R}}=1.45$ could reach $13\sigma$, assuming that the statistical uncertainty in helical asymmetry measurement is reduced by a factor of $1/\sqrt{10}$, compared to the present analysis using $10^{6}$ spins.
@@
I acknowledgments
We thank anonymous referees for helpful comments that helped improve the original manuscript. We also thank Donghui Jeong, Eiichiro Komatsu, Oliver Philcox, Shi-Fan Chen, and Tomomi Sunayama for providing valuable feedback on the manuscript. We also thank Academia Sinica Institute of Astronomy and Astrophysics (ASIAA) for hosting the workshop Large-scale Parity Violation Workshop during which this project was advanced. J.S. acknowledges the support by ASIAA. H.R.Y. is supported by National Science Foundation of China grant No. 12173030. T.O. acknowledges support from the Ministry of Science and Technology of Taiwan under grants No. MOST 111-2112-M- 001-061- and No. NSTC 112-2112-M-001-034- and the Career Development Award, Academia Sinica (AS-CDA-108-M02) for the period of 2019-2023. We thank ASIAA for providing computing resources (High Performance Computing Systems).